Two charged particles are placed at a distance 1.0cm.what is the minimum possible magnitude of the electric force acting on each charge?

Grade:12

1 Answers

prateek
7 years ago
now F=KQq/R^2
LET Q=q=1e.
F=K/R^2.
F=9X10^9 X 100 X 100 . (K=9X10^9  & R=1cm=0.01)
9x10^13 is the minimum force which can be attractive or repulsive depending upon the nature of charges.